OpenStudy (anonymous):

Let R be the region in the xy-plane between the graphs of y = ex and y = e-x from x = 0 to x = 2. a) Find the volume of the solid generated when R is revolved about the x-axis.

OpenStudy (amistre64):



OpenStudy (amistre64):

the revolution is simply an area of a circle; pi [f(x)]^2

OpenStudy (anonymous):

Can you please explain in depth

OpenStudy (amistre64):

\[pi\int_{0}^{2}(e^{2x}-e^{-2x})dx\] seems to be what it amounts to then

OpenStudy (amistre64):

in depth; which part?

OpenStudy (amistre64):



OpenStudy (anonymous):

I have to use big R and little r both squared because the shape is hollow right? so I take both the solid and hollow part as my big R squared and then just the hollow part as my little r squared?

OpenStudy (amistre64):

the radius of the circle whose area we want is determined by the function f(x) correct?

OpenStudy (amistre64):

yes, Big radius little radius

OpenStudy (anonymous):

Thats what you put in the integral right?

OpenStudy (amistre64):

yes\[R_1 ^2=(e^x)^2 = e^2x\]\[R_2 ^2=(e^{-x})^2=e^{-2x}\]

OpenStudy (anonymous):

okay thank you !

OpenStudy (amistre64):

if you get a negative result; dont worry, it just means you got your top and bottom function swapped; drop the negative and your answer will be good

OpenStudy (anonymous):

thanks :D

OpenStudy (amistre64):

youre welcome :)